# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v2

@@ 3.7.1 Autonomous flying performance
Obstacle avoidance reliability was evaluated based on the mission success. A mission was considered successful if the drone reached the end of the flight without major collisions. Minor touches to thin vegetation or branches were allowed if the drone was able to continue the mission after that. If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal. In addition to obstacle avoidance, the smoothness of the trajectory was assessed. The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
@@
$\text{ATE}_{\text{pos}}=\left(\frac{1}{N}\sum_{i=0}^{N-1}||\Delta\mathbf{p}_{i}||^{2}\right)^{\frac{1}{2}}$ (2)
@@
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
The loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets that could affect the system performance. Since there were no GCPs for forming the reference trajectories, only the ability to detect the loops inside dense natural forests was tested, and the accuracy of the loop-corrected trajectory was not evaluated. VINS-Fusion estimate with the loop detection algorithm was run for the collected data with the onboard computer of the drone, and the effect of a maximum number of features in pictures for loop detection capability was examined.
@@ 3.7.2 DBH estimation performance
The accuracy of the DBH estimation from the drone point cloud data was evaluated using TLS data as a reference. The TLS scannings were conducted in April 2020, and tree height, DBH, and other parameters were estimated from the dataset. The estimated error of the DBH measurement was less than 1 cm.
@@
[…]
Additionally standard deviation was calculated with following equation:
@@
The impact of tree growth between the collection of drone data and reference measurements (four growth seasons) was not taken into account. Variability in tree growth, influenced by factors such as species, locations, health, and age, is acknowledged as a source of error. Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates. Additionally, some reference trees fell between the measurements and were not included in the estimations.
@@ 4.1.1 Obstacle avoidance
The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest. Consistent with these expectations, the trajectories were smoother and the success rate was better in ’Evo-medium’ than in ’Evo-difficult’ test area. In ’Evo-medium’, all performed test flights were successful and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight. In ’Evo-difficult’, one of nine test flights failed. None of the test flights in ’Evo-difficult’ was performed with a smooth trajectory. However, despite the emergency stops during the flight, the system was still able to continue the mission after the stops, leading to a successful mission in eight of the nine test flights.
The reason for emergency stops in both environments was the late detection of thin and dry spruce branches. In the failed test flight in ’Evo-difficult’ the reason was also a collision to a similar thin low branch.
@@ 4.1.2 VIO
Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’ while Table 4 summarizes the VINS-Fusion estimation errors across all flights. The average $\text{ATE}_{\text{pos}}$ was 0.50 m in ’Evo-medium’ and 0.34 m in ’Evo-difficult’ with standard deviations of 0.06 m and 0.07 m, respectively, without online $t_{d}$ estimation. The VINS-Fusion estimates with online $t_{d}$ estimation were computed post-flight for the same trajectories, yielding an average $\text{ATE}_{\text{pos}}$ of 0.47 m in ’Evo-medium’ and 0.33 m in ’Evo-difficult’ with standard deviations of 0.02 m and 0.04 m, respectively.
@@
In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests. Figure 11 shows examples of detected loop closure image pairs from both paths. The number of detected loops can be increased by increasing the number of features per image in the VINS-Fusion configuration. By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path. By decreasing the minimum distance between features to 30 pixels, VINS-Fusion detected a loop closure between 82 image pairs in the shorter path and 63 image pairs in the longer path. However, increasing the number of features also increases the computational burden of the VIO, and the effects on the real-time performance during flying were not tested in this study.
@@ 4.2.2 Tree detection
Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees and the boundaries of the point clouds. I and II followed almost the same trajectory, while III and IV had similar trajectories. Dataset I trajectory was the smoothest, and it did not include any sharp turns. The other three trajectories had one or more sharper turns. For datasets I, II, III, and IV, the number of reference trees inside the point cloud boundary was 29, 29, 31, and 32, of which 23, 18, 19, and 20 were found from the drone-collected point clouds, respectively. The resulting completeness values were 79.31 %, 62.07 %, 61.29 % and 62.5 %, respectively. The total number of different trees found was 29 out of 33 (87,88 %) from dataset I-IV. However, when all the flights were processed simultaneously (combined) that resulted in only $3$ completeness, where only 12 of 36 trees were found. When utilizing 0, 1, 2, and 3 GCPs in the dataset I, completeness values were 72.41 %, 68.97 %, 75.86 %, and 79.31 %, respectively (Table 5).
@@
In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory despite these trees being well-reconstructed in the point cloud. Similarly, the number of well-reconstructed trees but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively. The three extra trees that were not identified in dataset I were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud. The datasets II through IV each contained seven poorly reconstructed trees. Many of these trees were located at a greater distance from the trajectory, while others were obscured from the view of the camera by intervening trees.
@@ 4.2.3 DBH estimation
Tree size had an impact on the DBH estimation accuracy (Table 7). For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved superior performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively. Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99 %-25.44 %) and -0.66 to 2.08 cm (-1.53 %-5.65 %), respectively.
@@
Table 7: Accuracy of the stem diameter breast height (DBH) in ’Evo-medium’ test site. Dataset is split in 3 categories: All, DBH $<$ 30 cm and DBH $>$ 30 cm trees. Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
Dataset Number of trees All $<$30 cm $>$30 cm RMSE (cm) All $<$30 cm $>$30 cm RMSE % All $<$30 cm $>$30 cm Bias (cm) All $<$30 cm $>$30 cm Bias % All $<$30 cm $>$30 cm Standard Deviation All $<$30 cm $>$30 cm
@@ 5.1 Autonomous flying
The performance of autonomous navigation was promising in both test forests. The obstacle avoidance success rate was 100 % in the ’Evo-medium’ forest and 87.5 % in the ’Evo-difficult’ forest. However, the smoothness of the trajectories suffered from the increasing forest density. While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’ at least one emergency stop was needed in all of the flights. However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory re-planning and finish the mission after the emergency stops. Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier. Nonetheless, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium-difficulty level boreal forests.
The accuracy of the position estimate by VINS-Fusion was satisfying for short flights performed in this study, but in the longer flights, the error would start to accumulate, making the position errors high. However, the standard deviation between the errors between the test flights was small, and the VINS-Fusion systematically estimated the flight paths shorter than in reality (as can be seen from Figure 9). This indicates that a notable part of the position error could be eliminated by improving the system calibration. The results indicate that the estimates by VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration. However, further development of the prototype should include implementing hardware-level synchronization between the camera and the IMU of the autopilot.
The position error can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion could detect loop closures even in a homogeneous forest environment. Increasing the number of features tracked by VINS-Fusion increased the number of detected loops. However, increasing the number of features also increases the computational burden of the system, potentially disrupting the real-time estimation. Finding the optimal value for the tracked features, as well as measuring the actual improvement of the loop closures to the position estimate, requires further studies.
@@ 5.2 Accuracy of forest measurements
The proposed system was assessed by comparing it to other relevant studies (Section 2.1, Table 1) considering the accuracy, system autonomy, weight, and complexity of the test environment.
The RMSEs reported in the relevant studies varied from 3.2 % to 61.41 %. The highest accuracy UAV SfM-MVS results were presented by Shimabuku, Konoshima, and Ota (2023) with RMSEs of 0.4 cm (3.2 %) and 0.7 cm (3.6 %) in two test forest sites. However, their test sites included only 7 and 18 reference trees, and the semi-autonomous commercial drone system needed manual control to avoid horizontal obstacles. Kuželka and Surovỳ (2018), He et al. (2025), and Krisanski, Taskhiri, and Turner (2020) had more reference trees and achieved similar DBH accuracy compared to that in this study, but the flights were piloted manually. Notably, the SfM-MVS approaches involved commercial lightweight drone systems, except He et al. (2025) where 360° camera was mounted as an additional payload to heavier LiDAR-based custom drone.
The most accurate under-canopy LiDAR-based DBH estimates were presented by Hyyppä et al. (2020a), achieving an RMSE of 0.6 cm (2.3 %) and bias of 0.34 cm (1.3 %) and 1.1 cm (3.5 %) and 0.12 cm (0.4 %) in two sparse forest plots. However, their system weighed around 10 kg and was piloted manually.
Autonomously operating under-canopy LiDAR-based drones were utilized in the studies by Muhojoki et al. (2024), Liang et al. (2024), and Cheng et al. (2024). While the commercial Hovermap drone system utilized in Muhojoki et al. (2024) achieved the lowest RMSE of these, the system was able to fly only in sparse and clean forests due to its size. The automated under-canopy UAV Lidar system by Liang et al. (2024) weighted 5 kg and provided an RMSE of 4.48 cm (21.37 %) and bias of 5.01 (14.91 %) at best. The autonomous system utilized in Cheng et al. (2024) weighed 4.2 kg and provided an RMSE of 3.8 cm (12.8 %) with the data collected in autonomous flight mode. The same LiDAR-based drone platform was used in studies by Prabhu et al. (2024) and He et al. (2025), but based on the articles, it seems that data collecting flights were performed in manual flight mode in these studies. Prabhu et al. (2024) reported RMSE of 1.67 cm (7.42 %) with LiDAR data, and He et al. (2025) reported median relative error of 11.5 % with photogrammetric point clouds from 360° camera data.
@@
In summary, the highest accuracy levels were obtained with the relatively heavy ULS systems operated in sparse forest environments or in environments with few reference trees. Miniaturized lightweight sensors were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al. (2024), Cheng et al. (2024), He et al. (2025), and Liang et al. (2024), though the LiDAR-based systems (Prabhu et al. 2024; Cheng et al. 2024; He et al. 2025; Liang et al. 2024) were still considerably heavier and larger than the camera-based drone systems. From these, only studies by Shimabuku, Konoshima, and Ota (2023), Cheng et al. (2024), and Liang et al. (2024) reported flights with a high level of automation, and LiDAR-based flights in Cheng et al. (2024) and Liang et al. (2024) were the only ones where the systems were able to fly without commands from the pilot during the flight. Considering the complexity of the environments, the density of the forest in Liang et al. (2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha). The natural-like subtropical forest in Shimabuku, Konoshima, and Ota (2023) had seven trees in 5 m $\times$ 5 m area, leading to a density of 2000 trees/ha, but the test plot had minimal size. In Cheng et al. (2024), the forest density was not reported.
@@
The approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a similar level of estimation accuracy to most of the manually piloted UAV SfM-MVS systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al. 2025). Compared to the studies with LiDAR-based autonomous UAV collection, the accuracy was higher than in Liang et al. (2024) and similar than in Cheng et al. (2024) even without any GCPs. However, the system utilized in Cheng et al. (2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al. 2024; Prabhu et al. 2024), suggesting that their LiDAR-based system has the potential for higher estimation accuracy by optimizing the autonomous flight trajectories.
Although the estimation accuracy obtained in this study was in line with most of the relevant studies even without any GCPs, the reconstruction quality (completeness and RMSE) can be further improved with GCPs.
Based on the previous, it can be concluded that camera-based autonomous miniaturized drones hold the potential to become a practical tool in forest parameter estimation.The potential estimation accuracy with low-cost stereo-cameras is near to the quality of small LiDARs suitable for miniaturized drones, but the weight and the price of the sensor are considerably lower. This allows the building of smaller platforms capable of flying in denser forests, as well as reduces the cost of the system.

### v3

@@ 3.7.1 Autonomous flying performance
The reliability of obstacle avoidance was evaluated on the basis of mission success. A mission was considered successful if the drone reached the end of the flight without major collisions. Minor touches to thin vegetation or branches were allowed if the drone was able to continue the mission after that. If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal. In addition to obstacle avoidance, the smoothness of the trajectory was assessed. The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
@@
[…]
@@
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to a constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of the online estimation.
The loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets that could affect the system performance. Since there were no GCPs to form the reference trajectories, only the ability to detect loops inside dense natural forests was tested and the accuracy of the loop-corrected trajectory was not evaluated. VINS-Fusion estimate with the loop detection algorithm was run for the collected data with the onboard computer of the drone, and the effect of a maximum number of features in pictures for loop detection capability was examined.
@@ 3.7.2 DBH estimation performance
The accuracy of the DBH estimation from the drone point cloud data was evaluated using TLS data as a reference. The TLS scans were collected in April 2020, and tree height, DBH, and other parameters were estimated from the dataset. The estimated error of the DBH measurement was less than 1 cm.
@@
$\text{Relative RMSE (\%)}=\frac{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)^{2}}}{\bar{d}_{\text{reference}}}\times 100$ (5)
$\text{Relative Bias (\%)}=\frac{\frac{1}{n}\sum_{i=1}^{n}\left(d_{\text{estimated},i}-d_{\text{reference},i}\right)}{\bar{d}_{\text{reference}}}\times 100$ (6)
Additionally, standard deviation was calculated with the following equation:
@@
The impact of tree growth between the collection of drone data and reference measurements (four growth seasons) was not taken into account. Variability in tree growth, influenced by factors such as species, locations, health, and age, is acknowledged as a source of error. Considering that the tree growth was relatively constant in the test area, it formed a major component of the bias in the estimates. Additionally, some reference trees fell between the measurements and were not included in the estimations.
@@ 4.1.1 Obstacle avoidance
The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest. Consistent with these expectations, the trajectories were smoother and the success rate was better in ’Evo-medium’ than in ’Evo-difficult’ test area. In ’Evo-medium’, all performed test flights were successful, and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight. In ’Evo-difficult’, one of nine test flights failed. None of the test flights in ’Evo-difficult’ was performed with a smooth trajectory. However, despite emergency stops during the flight, the system was still able to continue the mission after stops, leading to a successful mission on eight of the nine test flights.
The reason for emergency stops in both environments was the late detection of thin and dry spruce branches. In the failed test flight in ’Evo-difficult’, the reason was also a collision with a similar thin low branch.
@@ 4.1.2 VIO
Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’, while Table 4 summarizes the VINS-Fusion estimation errors across all flights. The average $\text{ATE}_{\text{pos}}$ was 0.50 m in ’Evo-medium’ and 0.34 m in ’Evo-difficult’ with standard deviations of 0.06 m and 0.07 m, respectively, without online $t_{d}$ estimation. The VINS-Fusion estimates with online $t_{d}$ estimation were computed post-flight for the same trajectories, yielding an average $\text{ATE}_{\text{pos}}$ of 0.47 m in ’Evo-medium’ and 0.33 m in ’Evo-difficult’ with standard deviations of 0.02 m and 0.04 m, respectively.
@@
In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests. Figure 11 shows examples of pairs of detected loop closure images from both paths. The number of detected loops can be increased by increasing the number of features per image in the VINS-Fusion configuration. By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path. By decreasing the minimum distance between features to 30 pixels, VINS-Fusion detected a loop closure between 82 image pairs in the shorter path and 63 image pairs in the longer path. However, increasing the number of features also increases the computational burden of the VIO, and the effects on the real-time performance during flying were not tested in this study.
@@ 4.2.2 Tree detection
Figure 12 shows post-processed trajectories for dataset I-IV, as well as reference and detected trees and the boundaries of the point clouds. I and II followed almost the same trajectory, while III and IV had similar trajectories. The dataset I trajectory was the smoothest, and it did not include any sharp turns. The other three trajectories had one or more sharper turns. For datasets I, II, III, and IV, the number of reference trees inside the point cloud boundary was 29, 29, 31, and 32, of which 23, 18, 19, and 20 were found from the drone-collected point clouds, respectively. The resulting completeness values were 79.31 %, 62.07 %, 61.29 % and 62.5 %, respectively. The total number of different trees found was 29 out of 33 (87,88 %) from dataset I-IV. However, when all flights were processed simultaneously (combined), that resulted in only 33.33 % completeness, where only 12 of 36 trees were found. When utilizing 0, 1, 2, and 3 GCPs in the dataset I, completeness values were 72.41 %, 68.97 %, 75.86 %, and 79.31 %, respectively (Table 7).
@@
In dataset I, the stem detection algorithm failed to detect one large tree and two small trees located near the flight trajectory, despite these trees being well reconstructed in the point cloud. Similarly, the number of trees well reconstructed but not detected by the algorithm was 4, 5, and 5, in datasets II, III, and IV, respectively. The three extra trees that were not identified in dataset I were situated further from the trajectory and exhibited poor reconstruction quality in the point cloud. The datasets II through IV each contained seven poorly reconstructed trees. Many of these trees were located at a greater distance from the trajectory, while others were obscured from the view of the camera by intervening trees.
@@ 4.2.3 DBH estimation
Tree size had an impact on the accuracy of DBH estimation (Table 7). For small and medium-sized trees with a DBH smaller than 30 cm, DBH estimations achieved good performance in all datasets, with RMSE and bias varying from 1.16 to 2.75 cm (5.74 %-12.47 %) and -0.01 to 0.62 cm (-0.06 %-3.14 %), respectively. Significantly poorer performance was obtained for trees with a DBH greater than 30 cm, with RMSE and bias ranging from 4.04 to 11.08 cm (8.99 %-25.44 %) and -0.66 to 2.08 cm (-1.53 %-5.65 %), respectively.
@@
Table 7: Accuracy of the stem diameter breast height (DBH) in ’Evo-medium’ test site. Dataset is split in 3 categories: All, DBH << 30 cm and DBH >> 30 cm trees. Datasets: I: Flight 1, II: Flight 2, III: Flight 3, IV: Flight 4, Combined: Flights I-IV merged, and for Dataset I: I0GCP: with no GCPs, I1GCP: with 1 GCP, I2GCP: with 2 GCPs, I3GCP with 3 GCPs.
Dataset Number of trees All <<30 cm >>30 cm RMSE (cm) All <<30 cm >>30 cm RMSE % All <<30 cm >>30 cm Bias (cm) All <<30 cm >>30 cm Bias % All <<30 cm >>30 cm Standard Deviation All <<30 cm >>30 cm
@@ 5.1 Autonomous flying
The performance of autonomous navigation was promising in both test forests. The obstacle avoidance success rate was 100 % in the ’Evo-medium’ forest and 87.5 % in the ’Evo-difficult’ forest. However, the smoothness of the trajectories suffered from the increase in forest density. While in ’Evo-medium’ 71.4 % of the flights had a smooth trajectory, in ’Evo-difficult’, at least one emergency stop was needed in all of the flights. However, apart from the one failed flight in ’Evo-difficult’, the system was able to do a successful trajectory replanning and finish the mission after the emergency stops. Further consideration should be given to finding the optimal sensor for detecting thin and dry branches earlier. However, the current prototype already allows reliable and safe autonomous obstacle avoidance in easy and medium difficulty level boreal forests.
The accuracy of the position estimate of VINS-Fusion was satisfactory for the short flights performed in this study, but during longer flights, the error would start to accumulate, resulting in high position errors. However, the standard deviation between the errors of the test flights was low and VINS-Fusion systematically estimated that the flight paths were shorter than in reality (as can be seen in Figure 9). This indicates that a notable part of the position error could be eliminated by improving the system calibration. The results indicate that the estimates of VINS-Fusion were more accurate with online $t_{d}$ estimation than when the $t_{d}$ value was fixed to the one obtained from the calibration. However, further development of the prototype should include implementation of hardware-level synchronization between the camera and the IMU of the autopilot.
Position errors can also be reduced by optimizing the estimate with loop closures, and the results from the walking tests show that VINS-Fusion can detect loop closures even in a homogeneous forest environment. Increasing the number of features tracked by VINS-Fusion increased the number of detected loops. However, increasing the number of features also increases the computational burden of the system, potentially disrupting real-time estimation. Finding the optimal value for the tracked features, as well as measuring the actual improvement of the loop closures to the position estimate, requires further studies.
@@ 5.2 Accuracy of forest measurements
The proposed methodology and the accuracy of the DBH estimates obtained were evaluated by comparing them with other relevant studies (Section 2.1, Table 1) considering the precision, autonomy of the system, weight, and complexity of the test environment.
The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems Hyyppä et al. (2020a, c); Muhojoki et al. (2024) operated in sparse forest environments or in environments with few reference trees Shimabuku, Konoshima, and Ota (2023). The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (Kuželka and Surovỳ 2018; He et al. 2025; Krisanski, Taskhiri, and Turner 2020). Furthermore, other studies based on SfM (Shimabuku, Konoshima, and Ota 2023; Kuželka and Surovỳ 2018; He et al. 2025; Krisanski, Taskhiri, and Turner 2020) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
@@
Considering the size of the systems, miniaturized lightweight systems were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al. (2024), Cheng et al. (2024), He et al. (2025), and Liang et al. (2024), though the LiDAR-based custom drone systems (Prabhu et al. 2024; Cheng et al. 2024; He et al. 2025; Liang et al. 2024) were still considerably heavier and larger than the camera-based drone systems. From these, LiDAR-based flights in Cheng et al. (2024) and Liang et al. (2024) were the only ones where the systems were reported to fly without commands from the pilot during the flight. Considering the complexity of the environments, the density of the forest in Liang et al. (2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha). In Cheng et al. (2024), the density of the forest was not reported.
@@
In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al. 2025). In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except He et al. (2025) where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported. Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in Liang et al. (2024) and close to that in Cheng et al. (2024) even without any GCPs. However, the system utilized in Cheng et al. (2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al. 2024; Prabhu et al. 2024), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
Based on the results, the potential estimation accuracy with a low-cost stereo camera is considerably lower than that with heavy and expensive survey LiDARs but close to the quality of smaller LiDARs suitable for miniaturized drones capable of flying inside dense forests. However, the weight and price of the sensor are considerably lower, allowing the building of smaller platforms capable of flying in even denser forests, as well as reducing the cost of the system. The results show that GCPs can be used to improve the estimation quality (completeness and RMSE). However, the increase in estimation error was relatively low when GCPs were not used in postprocessing (RMSE from 12.98 % to 14.66 %) suggesting that the approach has the potential to work even without additional GCPs.
Based on previous findings, it can be concluded that camera-based autonomous miniaturized drones have the potential to become a practical and cost-effective tool for forest parameter estimation in the future. However, further studies are required to increase the completeness of the reconstruction and the reconstruction coverage of the trees to decrease the RMSE of the estimation, especially for trees with large diameters.